OpenStudy (anonymous):

Evaluate the expression for the given variable(s). −2(4h+6)−6+h for h=−5

OpenStudy (anonymous):

\[-2(4h+6)-6+h\]h = -5 Alright, so, if h = -5, plug in the number into the equation.\[-2(4(-5)+6)-6+(-5) = ?\]Now you use PEMDAS -2(-20+6)-11 = ? -2(-16)-11 = ? 32 - 11 = ?